# Signal-noise separation using unsupervised reservoir computing

### v1

@@ I INTRODUCTION
Recovering the signal using data contaminated by noise is a long-standing challenge in signal and image processing Attias and Schreiner (1998); Moulines, Cardoso, and Gassiat (1997); Kundur and Hatzinakos (1996); Campisi and Egiazarian (2017). Estimation of the noise distribution is a key factor in determining the effectiveness of noise removal algorithms. In general, such algorithms are designed to work well for a specific distribution of noise. For example, algorithms that are designed to remove Gaussian noise will not perform as well for non-Gaussian noise. When handling denoising without prior knowledge of the noise distribution, it is often necessary to try a variety of algorithms such as low-pass filtering, wavelet-based smoothing and nonlinear adaptive filtering, and see which one works best Borodinov et al. (2019); Li et al. (2019). However, those trials may not be able to remove the noise effectively and often introduce artifacts to the recovered signal.
The signal-noise separation task becomes even more challenging if there is no reliable noise model. As the methods for dealing with multiplicative noise significantly differ from those employed for additive noise. Identifying whether the signal is corrupted by additive noise, multiplicative noise, or a combination of both is crucial for accurate signal analysis Coulon, Tourneret, and Swami (2000); Uhm et al. (1991).
@@ IV Reservoir Computing Predictor
The predictor $P$ in the process outlined in the previous section can be implemented with any common machine learning methods. Reservoir computing (RC) is one of the popular choices to deal with time series due to its simple architecture and dynamic nature Schrauwen, Verstraeten, and Van Campenhout (2007). RC comprises two main components: (i) a “reservoir” is a fixed nonlinear recurrent network and (ii) a “readout” is a trainable linear output layer.

### v2

@@ I INTRODUCTION
Recovering the signal using data contaminated by noise is a long-standing challenge in signal and image processing Attias and Schreiner (1998); Moulines et al. (1997); Kundur and Hatzinakos (1996); Campisi and Egiazarian (2017). Estimation of the noise distribution is a key factor in determining the effectiveness of noise removal algorithms. In general, such algorithms are designed to work well for a specific distribution of noise. For example, algorithms that are designed to remove Gaussian noise will not perform as well for non-Gaussian noise. When handling denoising without prior knowledge of the noise distribution, it is often necessary to try a variety of algorithms such as low-pass filtering, wavelet-based smoothing and nonlinear adaptive filtering, and see which one works best Borodinov et al. (2019); Li et al. (2019). However, those trials may not be able to remove the noise effectively and often introduce artifacts to the recovered signal.
The signal-noise separation task becomes even more challenging if there is no reliable noise model. As the methods for dealing with multiplicative noise significantly differ from those employed for additive noise. Identifying whether the signal is corrupted by additive noise, multiplicative noise, or a combination of both is crucial for accurate signal analysis Coulon et al. (2000); Uhm et al. (1991).
@@ IV Reservoir Computing Predictor
The predictor $P$ in the process outlined in the previous section can be implemented with any common machine learning methods. Reservoir computing (RC) is one of the popular choices to deal with time series due to its simple architecture and dynamic nature Schrauwen et al. (2007). RC comprises two main components: (i) a “reservoir” is a fixed nonlinear recurrent network and (ii) a “readout” is a trainable linear output layer.